### v3

# SoftVQ-VAE: Efficient 1-Dimensional Continuous Tokenizer

Abstract
Efficient image tokenization with high compression ratios remains a critical challenge for training generative models. We present SoftVQ-VAE, a continuous image tokenizer that leverages soft categorical posteriors to aggregate multiple codewords into each latent token, substantially increasing the representation capacity of the latent space. When applied to Transformer-based architectures, our approach compresses 256$\times$256 and 512$\times$512 images using as few as 32 or 64 1-dimensional tokens. Not only does SoftVQ-VAE show consistent and high-quality reconstruction, more importantly, it also achieves state-of-the-art and significantly faster image generation results across different denoising-based generative models. Remarkably, SoftVQ-VAE improves inference throughput by up to 18x for generating 256$\times$256 images and 55x for 512$\times$512 images while achieving competitive FID scores of 1.78 and 2.21 for SiT-XL. It also improves the training efficiency of the generative models by reducing the number of training iterations by 2.3x while maintaining comparable performance. With its fully-differentiable design and semantic-rich latent space, our experiment demonstrates that SoftVQ-VAE achieves efficient tokenization without compromising generation quality, paving the way for more efficient generative models. Code and model are released  [^1] https://github.com/Hhhhhhao/continuous_tokenizer..
Figure 1: ImageNet-1K 256$\times$256 and 512$\times$512 generation results of generative models trained on SoftVQ-VAE with 32 and 64 tokens.
1 Introduction
Denoising-based generative modeling has witnessed remarkable progress with recent advances, such as Diffusion Transformers (DiT) [80], Scalable Interpolant Transformers (SiT) [72], and Masked Auto-Regressive models with diffusion loss (MAR) [60], to name a few. Denosing-based generative modeling not only presents impressive generation results on a wide range of modality, including natural language [32, 122], images [83, 16, 88, 69, 25], videos [42, 8, 6], and audios [81, 24], but also shows potential to unify the understanding and generation capabilities across modalities in multi-modal language models [123, 110, 107, 109].
A core component of denoising-based generative models is the tokenizer [51, 38, 104, 84, 23, 117, 56, 112], which compresses the raw data of each modality into a set of latent tokens in either a discrete or continuous latent space. The compact latent space therefore allows for more efficient and better generative modeling [88]. Among previous efforts, Kullback–Leibler Variational Auto-Encoders (KL-VAE) [51] and Vector Quantized Variational Auto-Encoders (VQ-VAE) [104, 84, 23] stand out as representatives of tokenizers which introduce continuous and discrete latent spaces, respectively. The former constrains the latent space with a Gaussian distribution using re-parametrization [51] trick, and the latter makes its latent space a categorical discrete distribution with a codebook of finite vocabulary which requires straight-through estimation [7].
Although both KL-VAE and VQ-VAE (and their variants) have been predominantly adopted in many generative models [88, 80, 10, 98, 123], they still present two major challenges restricting the efficiency and effectiveness of generative modeling: (1) the difficulty of achieving a higher compression ratio [13] and (2) the worse discriminative representations in their latent space than other self-supervised methods [35, 14, 116, 109]. The efficiency of downstream generative models, particularly Transformer-based architectures [105], is fundamentally constrained by their quadratic complexity to the number of latent tokens. Current image tokenizers [23, 88, 96] typically compress 256$\times$256 images to at least 256 tokens and 512$\times$512 images to at least 1024 tokens, creating a significant computational bottleneck for both training and inference of generative models [80, 72, 60]. Many efforts have been made to reduce the number of latent tokens on the generative model side, such as merging [80, 72, 62], pooling [58, 94, 61], and others [70, 63, 12]. More studies have recently emerged to fundamentally reduce the token number of the tokenizer [30, 115, 62, 11]. For example, TiTok [115] uses 128 tokens, achieving generation results comparable to 256 tokens by adding one extra decoder. DC-AE [11] compresses the initial 1024 tokens of 512$\times$512 with 256 tokens. However, further increasing the compression ratio results in a significant degradation in the quality of reconstruction and therefore the generation [11].
Beyond the compression challenge, the quality of the learned representations of the tokenizer poses another critical limitation. While latent representations are crucial for both understanding [31, 109] and generation tasks [30, 64, 116, 62], current tokenizers usually struggle to capture more discriminative features. Earlier methods attempted to advance the training objectives and recipe to improve the latent space of the tokenizer [23, 19, 34, 124, 14]. Recently, [30, 118, 113, 124, 109, 62, 116] has also explored the representation alignment of latent tokens. However, due to the lossy nature [91, 77] of the smoothness constraints in KL-VAE and the discrete quantization in VQ-VAE [38, 13, 126, 27], current tokenizers usually fall short in learning the semantics of the latent space [109].
To address the aforementioned challenges, we present SoftVQ-VAE, a simple modification to VQ-VAE that converts it from a discrete tokenizer into a continuous one with a high compression ratio and a semantic-rich latent space. Specifically, we propose using a soft categorical posterior in VAE with a learnable codebook. This straightforward modification introduces a crucial capability: instead of the conventional one-to-one mapping between tokens and codewords in VQ-VAE, our approach enables the adaptive aggregation of multiple codewords into each latent token, substantially increasing the representation capacity of the latent space. We apply SoftVQ-VAE to the Transformer auto-encoder architecture [111, 30, 58, 115, 62], and we succeed in using much fewer 1-dimensional latent tokens (i.e., 32 and 64) for both the reconstruction and subsequent generation tasks. Since SoftVQ-VAE is now fully differentiable, it not only simplifies the learning of the codebook, not requiring the codebook loss or commit loss as in VQ-VAE [104, 23], but also enables better representation learning through direct alignment with pre-trained semantic features using a simple cosine similarity objective. Our approach generalizes the K-Means [67] latent space of VQ-VAE to Soft K-Means [22] and further to Gaussian Mixture Models (GMM) [86], while maintaining compatibility with existing VQ-VAE techniques [45, 112, 55, 62]. We comprehensively demonstrate the efficiency and effectiveness of the proposed SoftVQ-VAE in image reconstruction and generation tasks, and show that:
  • SoftVQ-VAE presents more consistent, robust, and high-quality reconstruction and generation results with various latent token lengths ranging from 256 to 32.
  • SoftVQ-VAE allows diffusion-based DiT [73], flow-based SiT [72], and autoregressive-based MAR [60] to achieve state-of-the-art generation results with only 64 and 32 tokens on both the 256$\times$256, with an FID of 1.78, and 512$\times$512 on ImageNet benchmark, with an FID of 2.21.
  • SoftVQ-VAE improves the inference throughput by 18x and 10x with 32 and 64 tokens to generate 256$\times$256 images, and 55x with 64 tokens for generating 512$\times$512 images, with significantly lower GFLOPs for all models.
  • SoftVQ-VAE presents a more discriminative latent space through representation alignment. It also facilitates training efficiency by significantly improving training throughput and reduces training iterations by 2.3x.
2 Preliminary
We present an overview of the tokenizers, i.e., KL-VAE [51] and VQ-VAE [104], and denoising-based generative models, i.e., DiT [80], SiT [72], and MAR [60], in this section.
2.1 Image Tokenizer
The architecture of image tokenizer generally resembles auto-encoders, consisting of two main components: an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$, parameterized by $\phi$ and $\theta$, respectively. Given an input image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$, the encoder $\mathcal{E}$ maps the high-dimensional $\mathbf{x}$ to a lower-dimensional latent representation $\mathbf{z}=\mathcal{E}(\mathbf{x};\phi)$. The latent representations can have varying shapes depending on the encoder architecture. We generally view the latent representation as a set of latent tokens $\mathbf{z}=[\mathbf{z}^{[0]},\mathbf{z}^{[1]},\dots,\mathbf{z}^{[L]}]\in\mathbb{R}^{L\times D}$, where $L$ is the number of tokens and $D$ is the dimension of the latent representation. For basic auto-encoders (AE), the decoder takes latent tokens as input and reconstructs the original signal $\hat{\mathbf{x}}=\mathcal{D}(\mathbf{z};\theta)$ by minimizing the reconstruction loss with $\mathbf{x}$.
Variational auto-encoders (VAE) [51] extend the basic AE by introducing a probabilistic perspective on the latent space. Specifically, VAE approximates a posterior distribution $p(\mathbf{z}|\mathbf{x})$ with a learned distribution $q_{\phi}(\mathbf{z}|\mathbf{x})$ from the encoder, assuming that $\mathbf{x}$ is generated by the unobserved latent $\mathbf{z}$ with a prior distribution $p(\mathbf{z})$. The decoder instead takes a set of sampled latent tokens $\mathbf{z}\sim q_{\phi}(\mathbf{z}|\mathbf{x})$ from the posterior and aims to learn the marginal likelihood of data in a generative process by optimizing the evidence lower bound (ELBO): $\max_{\phi,\theta}\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\log p_{\theta}(\mathbf{x}|\mathbf{z})-\mathcal{L}_{\mathrm{kl}}(q_{\phi}(\mathbf{z}|\mathbf{x})|p(\mathbf{z}))$. To make the optimization tractable, assumptions have been made on the prior, leading to different variants of VAE.
KL-VAE [51, 38] parametrizes both the prior and posterior as Gaussians. The prior $p(\mathbf{z})$ is assumed to be the isotropic unit Gaussian $\mathcal{N}(0,1)$. The continuous latent code is parameterized with posterior mean $\boldsymbol{\mu}_{\phi}$ and variance $\boldsymbol{\sigma}^{2}_{\phi}$ predicted by the encoder using the “re-parametrization” trick with a noise variable $\boldsymbol{\varepsilon}$ from standard Gaussian:
$\begin{split}&\textit{posterior: }q_{\phi}(\mathbf{z}|\mathbf{x})=\mathcal{N}(\mathbf{z};\boldsymbol{\mu}_{\phi}(\mathbf{x}),\boldsymbol{\sigma}^{2}_{\phi}(\mathbf{x}))\\ &\textit{latent: }\mathbf{z}=\boldsymbol{\mu}_{\phi}(\mathbf{x})+\boldsymbol{\sigma}_{\phi}(\mathbf{x})\odot\mathbf{\varepsilon},\quad\epsilon\sim\mathcal{N}(0,\mathbf{I})\\ &\textit{kl: }\mathcal{L}_{\mathrm{kl}}=-\frac{1}{2}\left(1+\log\left(\boldsymbol{\sigma}_{\phi}^{2}(\mathbf{x})\right)-\boldsymbol{\mu}_{\phi}(\mathbf{x})^{2}-\boldsymbol{\sigma}_{\phi}^{2}(\mathbf{x})\right)\end{split}$
While KL-VAE is widely used in diffusion models, it is known to have the “posterior-collapse” issue [13]. In addition, the KL loss weight usually imposes a trade-off between the quality of reconstruction and the smoothness of its latent space [100], preventing high compression ratio and the learning of semantics in the latent space of KL-VAE.
VQ-VAE [104, 23], in contrast, generates latent code in a discrete space by taking the posterior as a $K$-way deterministic categorical distribution against a learnable codebook $\mathcal{C}=[\mathbf{c}^{[0]},\dots,\mathbf{c}^{[K]}]\in\mathbb{R}^{K\times D}$, with $\mathbf{c}$ as the codeword:
$\begin{split}&\textit{posterior: }q_{\phi}(\mathbf{z}=k|\mathbf{x})=\begin{cases}1&\text{if }k=\operatorname*{arg\,min}_{j}\|\hat{\mathbf{z}}-\mathbf{c}^{[j]}\|_{2}\\ 0&\text{otherwise}\end{cases}\\ &\textit{latent: }\mathbf{z}=\mathbf{c}^{[k]},\quad\text{where }k=\operatorname*{arg\,min}_{j}\|\hat{\mathbf{z}}-\mathbf{c}^{[j]}\|_{2}\\ &\textit{kl: }\mathcal{L}_{\mathrm{kl}}=\log K\\ \end{split}$
Due to the non-differentiable $\operatorname*{arg\,min}$, a “straight-through” trick [7] is adopted to approximate the gradient of the encoder output $\hat{\mathbf{z}}$ by directly copying gradient from the decoder input $\mathbf{z}$. Since the reconstruction objective does not impose direct gradients on the codebook $\mathcal{C}$, VQ-VAE additionally uses a codebook loss as $\|\text{sg}[\hat{\mathbf{z}}]-\mathbf{c}\|_{2}$ to move codewords toward the encoder output, along with a commit loss $\|\hat{\mathbf{z}}-\text{sg}[\mathbf{c}]\|_{2}$ to prevent arbitrary growth of codewords, where $\text{sg}[\cdot]$ denotes stop-gradient. The broken gradient hinders high compression ratio and latent space learning [43, 27].
2.2 Denoising-based Generative Models
Denoising-based generative models synthesize images by progressively transforming Gaussian noise $\boldsymbol{\varepsilon}\sim\mathcal{N}(0,\mathbf{I})$ into tokenizer latent codes $\mathbf{z}$ through a forward process [92, 41]:
$\mathbf{z}_{t}=\alpha_{t}\mathbf{z}+\sigma_{t}\varepsilon,$ (3)
where $\alpha_{t}$ and $\sigma_{t}$ are a decreasing and increasing function of $t$, and $t\in[0,T]$ [50]. In this paper, we adopt three approaches to implementing this denoising process as follows.
DiT [80] is a diffusion-based model with a transformer architecture. It formulates the process through a forward-time stochastic differential equation (SDE), where $\mathbf{z}_{t}$ converges to $\mathcal{N}(0,\mathbf{I})$ as $t\rightarrow T$. Generation occurs via a reverse-time SDE [2], with the model trained to predict noise $\boldsymbol{\varepsilon}$ at randomly sampled timesteps. Due to the long token length of the tokenizer [96] used in DiT, it has several variants to merge tokens at the input of the transformer backbone.
SiT [72] instead uses stochastic interpolants [1] and performs denoising via probability flow ordinary differential equation (PF ODE). The model learns the ODE’s velocity field [65] by minimizing mean-squared-error (MSE) loss. This velocity field defines a score function for generation, analogous to DiT’s reverse-time SDE.
MAR [60] combines diffusion with autoregressive generation. Unlike DiT and SiT’s parallel denoising of all latent tokens, MAR adopts an encoder-decoder transformer architecture [35] that progressively denoises tokens in a “next-set” autoregressive manner [99], starting from masked tokens, similar to MaskGIT [10] and MAGE [59].
3 Method
In this section, we present SoftVQ-VAE, a novel 1D continuous tokenizer with a high compression ratio and a semantic-rich latent space. We first introduce the architecture and the formulation of SoftVQ-VAE. Then, we show that SoftVQ-VAE can learn semantics easily by aligning its latent tokens with pre-trained features via its fully-differentiable property.
3.1 Architecture
Figure 2: Illustration of SoftVQ-VAE. Left: Transformer encoder-decoder architecture with image tokens, arbitrary length of latent tokens, and mask tokens. Right top: fully-differentiable SoftVQ illustration. Right bottom: latent space representation alignment.
We leverage Vision Transformer (ViT) [21, 111] as the architecture for the encoder $\mathcal{E}$ and decoder $\mathcal{D}$ of SoftVQ-VAE. Similarly to Yu et al. [115], instead of using the image features as latent, we initialize a set of extra 1D learnable tokens and use these tokens for reconstruction and subsequent generation. With the self-attention mechanism, it allows the learnable tokens to adaptively aggregate different image tokens to obtain the latent tokens of SoftVQ, detailed in Sec. 3.2. Specifically, for the encoder transformer with a patch size of $P$, an image $\mathbf{x}\in\mathbb{R}^{H\times W\times 3}$ is patchified into a sequence of image tokens $\mathbf{x}_{p}\in\mathbb{R}^{N\times D}$, where $N=\ HW/P^{2}$ and $D$ represent the token dimension size. We then concatenate a set of learnable tokens $\mathbf{z}_{l}\in\mathbb{R}^{L\times D}$ of length $L$ along with the image tokens as the input to the encoder, and retain only the output corresponding to the learnable latent tokens as the output of our encoder: $\hat{\mathbf{z}}=\mathcal{E}([\mathbf{x}_{p};\mathbf{z}_{l}];\phi)$.
For the decoder, we reconstruct images from a sequence of learnable mask tokens $\mathbf{m}\in\mathbb{R}^{N\times D}$ [5, 35], concatenated with latent tokens: $\hat{\mathbf{x}}=\mathcal{D}([\mathbf{m};\mathbf{z}];\theta)$. We use a linear layer at the end of the decoder to regress the pixel values from the mask tokens. For the image tokens at the encoder input and mask tokens at the decoder input, we apply 2D absolute position embeddings and 1D absolute position embedding for the latent tokens. Adopting more advanced position embedding techniques, such as RoPE [97, 36], may lead to better results and is left for future work. This design allows SoftVQ-VAE to use latent codes of arbitrary length, each adaptively associated with image tokens, for reconstruction and subsequent generative modeling, and model images of various resolutions using the same length of latent codes. An overview of the model architecture is shown in Fig. 2.
3.2 SoftVQ-VAE
As discussed above, the high compression ratio of both KL-VAE and VQ-VAE usually results in a significant degradation of quality in reconstruction and latent space. To overcome these limitations, we propose SoftVQ-VAE, a simple modification to VQ-VAE, which maintains the advantages of learning codewords to capture the data distribution while bridging it with more representation capacity as a continuous tokenizer. The core idea of SoftVQ-VAE is to allow each latent code to adaptively aggregate multiple codewords from the learnable codebook. Consequently, it imposes a soft categorical distribution on the posterior from the encoder:
posterior: $\begin{split}&\textit{posterior: }q_{\phi}(\mathbf{z}|\mathbf{x})=\operatorname*{Softmax}\left(-\frac{\|\hat{\mathbf{z}}-\mathcal{C}\|_{2}}{\tau}\right)\\ &\textit{latent: }\mathbf{z}=q_{\phi}(\mathbf{z}|\mathbf{x})\mathcal{C}\\ &\textit{kl: }\mathcal{L}_{\mathrm{kl}}=H(q_{\phi}(\mathbf{z}|\mathbf{x}))-H\left(\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}q_{\phi}(\mathbf{z}|\mathbf{x})\right)\\ \end{split}$
where $\tau$ is the temperature parameter controlling the sharpness of the softmax probability and is set to $0.07$ by ablation. The derivation of $\mathcal{L}_{kl}$ is shown in Sec. B.1.
Despite this simple modification, SoftVQ-VAE can significantly reduce the latent token length while maintaining a high quality of the reconstruction and latent space. More importantly, with a soft categorical distribution, SoftVQ-VAE is fully-differentiable, and thus the encoder and codebook can be optimized directly from the reconstruction loss (and other losses). It also allows for easier regularization of various forms on the latent space that drastically improves its quality, and less hyper-parameter tuning without codebook and commit loss [43] in the reduced training objectives.
Interpretation. The posterior and latent space of VQ-VAE can be interpreted with K-Means [67]. Hence, SoftVQ-VAE can be viewed as soft K-Means [22], with codewords as learnable prototypes in the latent space. This latent space can easily be extended to Gaussian Mixture Models (GMM) [86], which we refer to as GMMVQ-VAE. Instead of using fixed temperature, GMMVQ-VAE predicts data-dependent weights for the codeword prototypes, i.e., Gaussian means, to compute the soft categorical posterior: $q_{\phi}(\mathbf{z}|\mathbf{x})=\operatorname*{Softmax}(-\omega(\hat{\mathbf{z}})\|\hat{\mathbf{z}}-\mathcal{C}\|_{2})$. Although prior work has explored GMM-based VAEs [17, 47], our GMMVQ-VAE differs by learning discrete prototypes rather than continuous latent variables. While providing the additional benefit of interpretable Gaussian components in the latent space, GMMVQ achieves reconstruction quality and downstream generation performance similar to SoftVQ in our experiments, and thus we mainly adopt the simpler SoftVQ throughout this paper. More details are in Sec. B.2.
3.3 Representation Alignment of Latent Space
While the reconstruction quality of tokenizer is important, learning a high-quality latent space is more crucial for downstream denoising-based generative modeling [116]. Several recent approaches have explored aligning the latent code with pre-trained features through a contrastive loss [30, 109, 62] and initializing the codebook with pre-trained features [124]. However, imposing effective regularization on the latent remains a challenge due to the discrete quantization of VQ-VAE and the Gaussian constraints in KL-VAE.
Thanks to the fully-differentiable property of SoftVQ-VAE, we can now directly impose regularization on the latent space. Inspired by the recent work of REPA [116], we propose aligning the representations of latent codes with pre-trained vision encoders to facilitate learning the latent space. Compared to REPA, which aligns the representation at intermediate layers of downstream generative models, our method of feature alignment in the tokenizer’s latent space is equivalent to performing REPA at the input space of generative models. Specifically, to align the latent code with the image tokens from a pre-trained vision encoder, we replicate each latent token by $N/L$ times:
$\mathbf{z}_{r}=[\underbrace{\mathbf{z}^{[0]},\ldots,\mathbf{z}^{[0]}}_{N/L\text{ times}},\underbrace{\mathbf{z}^{[1]},\ldots,\mathbf{z}^{[1]}}_{N/L\text{ times}},\ldots,\underbrace{\mathbf{z}^{[L]},\ldots,\mathbf{z}^{[L]}}_{N/L\text{ times}}],$ (5)
and encourage the token-wise similarity with image tokens. We employ a projector multilayer perceptron (MLP) on top of the latent tokens to match the pre-trained feature $\mathbf{y}_{*}$:
$\mathcal{L}_{\mathrm{align}}=\frac{1}{N}\sum_{n=1}^{N}\operatorname{sim}\left(\mathbf{y}_{*}^{[n]},\text{MLP}\left(\mathbf{z}_{r}^{([n]}\right)\right).$ (6)
This alignment ensures the latent space of SoftVQ-VAE captures semantically discriminative features that are beneficial for training the downstream generative models, even if it does not directly translate into improved reconstruction performance from the tokenizer. However, as we will show in Sec. 4.4, the generation quality of trained generative models often depends more on the semantic structure of the latent space than on the tokenizer’s reconstruction capabilities.
3.4 Final Training Objective
The training objective of SoftVQ-VAE combines the reconstruction loss, perceptual loss [54, 46, 20, 120] and adversarial loss [33, 44] as in VQ [23], and representation alignment:
$\mathcal{L}=\mathcal{L}_{\textrm{recon}}+\lambda_{1}\mathcal{L}_{\textrm{percep}}+\lambda_{2}\mathcal{L}_{\textrm{adv}}+\lambda_{3}\mathcal{L}_{\textrm{align}}+\lambda_{4}\mathcal{L}_{\textrm{KL}},$ (7)
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, and $\lambda_{4}$ are hyper-parameters. Neither codebook nor commit loss as in VQ is needed. The perceptual loss $\mathcal{L}_{\textrm{percep}}$ helps capture high-level perceptual features, $\mathcal{L}_{\textrm{adv}}$ encourages the decoder to generate realistic images by removing the artifacts from the reconstruction loss only, $\mathcal{L}_{\textrm{align}}$ guides the latent space to align with pre-trained features, and $\mathcal{L}_{\textrm{kl}}$ as the KL divergence term in the ELBO.
4 Experiments
In this section, we validate the efficiency and effectiveness of SoftVQ-VAE with extensive experiments.
Table 1: System-level comparison on ImageNet 256$\times$256 conditional generation. We compare with both diffusion-based models and auto-regressive models with different types of tokenizers. Grey denotes SoftVQ. indicates results with DPM-Solver and 50 steps.
Gen. Model Tok. Model # Tokens $\downarrow$ Tok. rFID $\downarrow$ Gflops $\downarrow$ Throughput (imgs/sec) $\uparrow$ w/o CFG w/ CFG
gFID $\downarrow$ IS $\uparrow$ gFID $\downarrow$ IS $\uparrow$
Auto-regressive Models
Taming-Trans. [23] VQ 256 7.94 - - 5.20 290.3 - -
RQ-Trans. [55] RQ 256 3.20 908.91 7.85 3.80 323.7 - -
MaskGIT [10] VQ 256 2.28 - - 6.18 182.1 - -
MAGE [59] VQ 256 - - - 6.93 195.8 - -
LlamaGen-3B [98] VQ 256 2.19 781.56 2.90 - - 3.06 279.7
TiTok-S-128 [115] VQ 128 1.61 33.03 6.50 - - 1.97 281.8
MAR-H [60] KL 256 1.22 145.08 0.12 2.35 227.8 1.55 303.7
SoftVQ-L 32 0.61 67.93 2.08 3.83 211.2 2.54 273.6
MAR-H SoftVQ-BL 64 0.65 86.55 0.89 2.81 218.3 1.93 289.4
Diffusion-based Models
LDM-4 [88] KL 4096 0.27 157.92 0.37 10.56 103.5 3.60 247.7
U-ViT-H/2 [4] KL 1024 0.62 128.89 0.98 - - 2.29 263.9
MDTv2-XL/2 [29] 125.43 0.59 5.06 155.6 1.58 314.7
DiT-XL/2 [80] 80.73 0.51 9.62 121.5 2.27 278.2
SiT-XL/2 [72] 81.92 0.54 8.30 131.7 2.06 270.3
+ REPA [116] 5.90 157.8 1.42 305.7
SoftVQ-B 0.89 8.94 9.83 113.8 3.91 264.2
SoftVQ-BL 0.68 8.81 9.22 115.8 3.78 266.7
DiT-XL SoftVQ-L 32 0.74 14.52 8.74 9.07 117.2 3.69 270.4
SoftVQ-B 0.88 4.70 6.62 129.2 3.29 262.5
SoftVQ-BL 0.65 4.59 6.53 131.9 3.11 268.3
DiT-XL SoftVQ-L 64 0.61 28.81 4.51 5.83 141.3 2.93 268.5
SoftVQ-B 0.89 10.28 7.99 129.3 2.51 301.3
SoftVQ-BL 0.68 10.12 7.67 133.9 2.44 308.8
SiT-XL SoftVQ-L 32 0.74 14.73 10.11 7.59 137.4 2.44 310.6
SoftVQ-B 0.88 5.51 5.98 138.0 1.78 279.0
SoftVQ-BL 0.65 5.42 5.80 143.5 1.88 287.9
SiT-XL SoftVQ-L 64 0.61 29.23 5.39 5.35 151.2 1.86 293.6
4.1 Experiments Setup
Implementation Details of SoftVQ-VAE. We use the LlamaGen codebase [98] to build our SoftVQ-VAE. We instantiate 4 configurations: SoftVQ-S, SoftVQ-B, SoftVQ-BL, SoftVQ-L, with a total of 45M, 173M, 391M, and 608M parameters, respectively. Each configuration has variants of latent codes $L=64$ and $L=32$. For the representation alignment, we choose DINOv2 [79] for the pre-trained features, similarly as Yu et al. [116] and Li et al. [62]. To learn a better semantic latent space, we initialize our encoder from the pre-trained DINOv2 weights. Alignment with other pre-trained features is explored in Sec. 4.4. We train the tokenizers on ImageNet [15] of resolution 256$\times$256 and 512$\times$512 for 250K iterations. Similarly to Tian et al. [99] and Li et al. [62], we adopted the same frozen DINO-S [9, 79] discriminator with a similar architecture to StyleGAN [48, 49]. We use DiffAug [121], consistency regularization [119], and LeCAM regularization [101] for discriminator training as in [99]. For the training objective, we set $\lambda_{1}=1.0$, $\lambda_{1}=0.2$, $\lambda_{3}=0.1$, and $\lambda_{4}=0.01$, following previous common practice. More training details are shown in Sec. C.1.
[…]
Evaluation. We use reconstruction Frechet Inception Distance (rFID) [37] on ImageNet validation set to evaluate the tokenizer. To evaluate the performance of generation tasks, we report generation FID (gFID), Inception Score (IS) [90], Precision and Recall [52] (in Sec. D.5), with and without classifier-free guidance (CFG) [40]. We measure the efficiency of the generative models by GLOPs of the model’s forward pass on the latent codes of tokenizers, and training and inference throughput for the models using floating point 32 and a batch size of $64$ on a single AMD MI250.
4.2 Main Results
We present the main reconstruction and generation results on the ImageNet benchmark of resolution 256$\times$256 and 512$\times$512 in Tab. 1 and Tab. 2, respectively. We show that SoftVQ-VAE achieves reconstruction and generation performance (with generative models) comparable to leading systems with 32 and 64 tokens, while presenting a significant improvement in both training and inference efficiency.
Consistent high-quality reconstruction. SoftVQ variants achieve remarkable reconstruction quality with a high compression ratio. With only 32 and 64 tokens, our models maintain consistent rFID scores, i.e., 0.61-0.89 on the 256$\times$256 and 0.64-0.71 on 512$\times$512 benchmark, substantially outperforming VQ tokenizers with 256 tokens used in autoregressive models such as MaskGIT [10] and TiTok [115]. Our results are also comparable to KL tokenizers used in previous denoising-based models, using 32x fewer tokens. More discussion of different tokenizers is given in Sec. 4.3.
Table 2: System-level comparison on ImageNet 512$\times$512 conditional generation. We compare with both diffusion-based models and auto-regressive models with different types of tokenizers. Grey denotes SoftVQ. indicates results with DPM-Solver and 50 steps.
Gen. Model Tok. Model # Tokens $\downarrow$ Tok. rFID $\downarrow$ Gflops $\downarrow$ Throughput (imgs/sec) $\uparrow$ w/o CFG w/ CFG
gFID $\downarrow$ IS $\uparrow$ gFID $\downarrow$ IS $\uparrow$
Generative Adversarial Models
BigGAN [10] - - - - - - - 8.43 177.9
StyleGAN-XL [48] - - - - - - - 2.41 267.7
Auto-regressive Models
MaskGIT [10] VQ 1024 1.97 - - 7.32 156.0 - -
TiTok-B [115] VQ 128 1.52 - - - - 2.13 261.2
MAR-H SoftVQ-BL 64 0.71 86.55 1.50 8.21 152.9 3.42 261.8
Diffusion-based Models
ADM [16] - - - - - 23.24 58.06 3.85 221.7
U-ViT-H/4 [4] KL 4096 0.62 128.92 0.58 - - 4.05 263.8
DiT-XL/2 [80] 373.34 0.10 9.62 121.5 3.04 240.8
SiT-XL/2 [72] 373.32 0.10 - - 2.62 252.2
DiT-XL [11] AE 256 0.22 80.75 1.02 9.56 - 2.84 -
UViT-H [11] 128.90 6.14 9.83 - 2.53 -
UViT-H [11] AE 64 0.22 32.99 15.23 12.26 - 2.66 -
UViT-2B [11] 104.18 9.44 6.50 - 2.25 -
SiT-XL SoftVQ-BL 64 0.71 29.23 5.38 7.96 133.9 2.21 290.5
Generation performance comparable to SOTA. Notably, with only 32 tokens, the generation performance of DiT-XL and SiT-XL trained on all variants of the proposed SoftVQ significantly outperforms DiT-XL/2 and SiT-XL/2 with KL tokenizers of 1024 tokens, without using CFG. DiT-XL with SoftVQ-L of 32 tokens presents a 0.62 improvement on FID and SiT-XL with SoftVQ-L of 32 tokens shows a 0.71 improvement on FID, without using CFG. The improvement is further strengthened with 64 tokens. SiT-XL with SoftVQ-L of 64 tokens outperforms SiT-XL/2 with REPA [116], without using CFG. Applying CFG, SiT-XL with SoftVQ of 64 tokens achieves a performance comparable to the leading systems with the best FID as 1.78 on 256 and 2.21 on 512 benchmark. Using MAR-H [60], SoftVQ-BL and SoftVQ-L with 64 and 32 tokens present slightly worse results, compared to the KL tokenizer of 256 tokens, possibly due to the lowered learning rate. MAR-H and SiT-XL results on 512 benchmark with 32 tokens are included in Sec. D.2.
Efficiency. The system-level comparison reveals SoftVQ’s exceptional efficiency across multiple metrics. When integrated with MAR-H at 256$\times$256 resolution, SoftVQ reduces GFLOPs by  40% compared to standard KL tokenizer, i.e., 86.55 vs 145.08 GFLOPs with 64 tokens and further reducing to 67.93 GFLOPs with 32 tokens. SoftVQ enables DiT-XL and SiT-XL to achieve significantly lower computational costs i.e., 28.81 and 29.23 GFLOPs, respectively with 64 tokens, compared to their KL counterparts (80.73 GFLOPs with 1024 tokens). Importantly, these efficiency improvements come without compromising generation quality. SiT-XL trained on SoftVQ-BL with 64 tokens presents  55x throughput than the KL counterpart, while showing competitive IS and FID scores in all configurations. Not only do we require fewer training iterations to achieve a comparable performance to leading systems, i.e., 2.5M vs 7M steps of SiT-XL and DiT-XL, the high compression ratio of SoftVQ also improves the training throughput, reducing the time to train SiT-XL for 400K steps from 72 to 20 hours on 8 MI250.
4.3 Comparison of Tokenizers
We compare the proposed SoftVQ-VAE with the concurrent efficient image tokenizers, i.e., TiTok [115] and DC-AE [11]. To show the superiority of SoftVQ-VAE in achieving a high compression ratio, we additionally train VQ-VAE and AE using the small (S) configuration and the same training recipe of SoftVQ. To validate the generation performance, we train a SiT-L for 400K steps and report gFID and IS on 256 ImageNet without using CFG, as shown in Tab. 3.
Scalable performance. With a much smaller model size, i.e., 46M vs 390M, SoftVQ-S significantly outperforms both TiTok variants at 64 tokens, achieving better reconstruction quality with an rFID of 1.03 compared to 1.25, and better generation performance with a gFID of 11.24 and IS of 89.4 compared to the best TiTok gFID of 19.23 and IS of 61.8. Noteworthy is that the generation results SiT-L trained on SoftVQ with 32 tokens outperform those trained on KL tokenizer with 1024 tokens [72] by a large margin, i.e., 5.9 in gFID. Compared to DC-AE, SoftVQ-S demonstrates competitive scalability across token counts while maintaining significant generation quality even with fewer parameters.
Table 3: Comparison of tokenizers on class-conditional 256$\times$256 ImageNet. We report rFID, and gFID and IS of SiT-L without CFG.
Tokenizer # Params # Tokens Dim. rFID $\downarrow$ SiT-L
gFID $\downarrow$ IS $\uparrow$
KL [96] 676M 1024 4 0.62 18.79 72.0
TiTok-BL-KL [115] 389M 64 16 1.25 23.35 54.7
TiTok-BL-VQ [115] 390M 64 16 2.06 19.23 61.8
DC-AE-f32 [11] 323M 64 32 0.69 15.36 76.1
DC-AE-f54 [11] 677M 16 128 0.81 24.66 49.1
VQ-S 46M 256 8 1.45 19.02 58.6
128 16 2.61 18.51 62.4
64 32 4.04 20.07 54.5
32 64 10.97 21.57 47.6
AE-S 46M 256 8 1.15 22.11 58.3
128 16 1.33 22.31 58.5
64 32 1.64 25.53 54.3
32 64 2.01 28.18 45.5
SoftVQ-S 46M 256 8 0.80 9.21 93.6
128 16 0.92 10.12 85.8
64 32 1.03 11.24 89.4
32 64 1.24 12.89 79.5
Less Lossy Property. SoftVQ-S exhibits remarkably consistent performance across different compression ratios. When reducing tokens from 256 to 32, SoftVQ-S maintains a minimal degradation in rFID from 0.80 to 1.24, significantly outperforming both VQ-S, from 1.45 to 10.97, and AE-S, from 1.15 to 2.01. The benefits of the fully-differentiable property of SoftVQ along with representation alignment are particularly evident in generation quality, where SoftVQ-S achieves substantially better gFID of 9.21-12.89, and IS scores of 79.5-93.6, compared to baselines, showing its robustness in preserving information at high compression ratios and a better latent space for generative models.
4.4 Discussions on the Latent Space
We perform more analysis on the latent of SoftVQ here.
Alignment with different initialization and target models. As shown in Tab. 4, our results demonstrate the effectiveness of alignment across various pre-trained models. The encoder initialization and alignment with DINOv2-B [79] achieve superior performance with rFID 0.88 and IS 103.4, compared to using either component alone. Further improvements are observed with CLIP-B [82] and EVA-02-B [26] on reconstruction. We reveal that a better rFID does not necessarily translate to a better gFID. Instead, the latent space quality, reflected by linear probing accuracy, is more closely related to the performance of the generative models.
Linear probing of tokenizer and generative model. Linear probing results in Fig. 3 (details in Sec. C.3) reveal SoftVQ’s superior representation quality at both tokenizer and generative model. At the tokenizer level, the linear probing accuracy of SoftVQ variants (B, BL, L) consistently outperforms VQ-S and AE-S across all token counts, with the gaps widening at lower token counts, similarly observed by Yu et al. [115]. This advantage transfers to the generative model with the best linear probing accuracy on the intermediate features of SiT trained with SoftVQ.
Latent space visualization. UMAP visualizations on the latent space in Fig. 4 demonstrate SoftVQ’s ability to maintain structured and discriminative latent representations with minimal variance between encoder output ($\hat{\mathbf{z}}$) and decoder input ($\mathbf{z}$). Comparing Fig. 4 (d) and (c), we also observe the larger encoder can learn a more discriminative latent space, showing the superiority of SoftVQ’s representation learning.
4.5 Ablation Studies
We present a series of ablation studies in Sec. D.1, including SoftVQ variants with product quantization (PQ) [45], residual quantization (RQ) [55], and GMMVQ, codebook size, latent space size, and the temperature of SoftVQ. The results show the compatibility of SoftVQ with PQ and RQ, with slightly improved performance. GMMVQ presents results comparable to SoftVQ. Moreover, SoftVQ shows robustness to other parameters, such as codebook size and softmax temperature. We found that while increasing the dimension of the latent space can result in a significant improvement in reconstruction, it leads to deterioration in generation performance, possibly due to the learning difficulty of generative models with larger dimensions [92].
Table 4: Representation alignment with different encoder initialization and the target alignment models using SoftVQ-B 64 tokens.
Enc. Init. Align. Target rFID $\downarrow$ L.P. $\uparrow$ SiT-L
gFID $\downarrow$ IS $\uparrow$
- - 0.98 5.42 20.33 89.4
- DINOv2-B [79] 0.93 41.08 10.96 101.9
DINOv2-B [79] - 0.73 11.87 17.20 91.5
DINOv2-B [79] DINOv2-B [79] 0.88 42.42 10.13 103.4
CLIP-B [82] CLIP-B [82] 0.75 47.12 9.89 109.3
EVA-02-B [26] EVA-02-B [26] 0.81 44.23 10.34 112.7
5 Related Work
Image tokenization has emerged as a fundamental technique to bridge various vision tasks. Traditional approaches using autoencoders [39, 106] laid the groundwork by compressing images into latent representations. This field subsequently diverged into two main branches: generation-focused and understanding-focused approaches. Generation-oriented methods, such as VAE [104, 84] and VQ-GAN [23, 85], emphasized learning latent spaces for detail-preserving compression. These were further refined through variants [55, 112, 73, 124] that enhanced generation quality. Parallel developments in understanding-focused approaches leveraged Large Language Models (LLMs) [105] to create semantic representations for tasks like classification [21] and detection [125]. Recent work [115, 62] has demonstrated the potential of unifying these approaches, with subsequent research [109, 34] exploring the convergence of generation and understanding capabilities using a single tokenizer.
Image generation have two popular paradigms: autoregressive and diffusion. Autoregressive models start from CNN-based architectures [103] and evolve to the transformer-based architectures [105, 114, 55, 66, 98] to enhance scalability. Latest works in visual autoregressive modeling, including VAR [99], MAR [60], and ImageFolder [62], have further enhanced generation efficiency. Diffusion models have transformed image generation since their introduction by Sohl-Dickstein et al. [93]. Key advancements by Nichol et al. [75], Dhariwal et al. [16], and Song et al. [95] have significantly improved model performance. A paradigm shift occurred with the adoption of latent space diffusion [102, 88], leveraging pre-trained image encoders as priors [104, 23] to enhance generation efficiency. Architectural innovations, particularly the integration of transformers [80], have further expanded the capabilities of these models. These advances have spawned numerous influential systems, from text-guided generation [76, 18] to high-fidelity synthesis [28, 89]. Commercial successes [83, 87, 74, 78] have also demonstrated the practical impact of diffusion models.
Figure 3: Linear probing accuracy of ImageNet-1K val. set on (a) latent tokens of tokenizer and (b) intermediate features (layer 20) of SiT (L for small and XL for others) trained on latents of tokenizer.
Figure 4: Visualization of $\hat{\mathbf{z}}$ (top), i.e., encoder output, and $\mathbf{z}$ (bottom), i.e., decoder input, of (a) VQ-S 64; (b) SoftVQ-S 64; (c) SoftVQ-S 32; (d) SoftVQ-L 32, trained with latent space alignment.
6 Conclusion
We presented SoftVQ-VAE, a continuous image tokenizer leveraging soft categorical posteriors to aggregate multiple codewords into each latent token. Our approach compresses images to just 32 or 64 tokens while maintaining high reconstruction quality and enabling state-of-the-art generation results with DiT, SiT, and MAR. The substantial improvements in both inference throughput (up to 55x with 64 tokens) and training efficiency, combined with the fully-differentiable design’s ability to learn semantic-rich representations, demonstrate SoftVQ-VAE’s potential as a foundation for efficient generative and vision-language models.
Supplementary Material
Appendix A Posterior of KL-VAE and VQ-VAE
In this section, we provide the detailed derivation of the KL-divergence of KL-VAE and VQ-VAE used in Eq. 1 and Eq. 2, respectively.
KL-VAE has the KL divergence of the posterior with a Gaussian prior:
$\mathcal{L}_{\mathrm{kl}}\left(q_{\phi}(\mathbf{z})\|p(\mathbf{z})\right).$ (8)
[…]
$=\frac{1}{2}\sum_{j=1}^{D}\left(1+\log\left(\left(\sigma_{j}\right)^{2}\right)-\left(\mu_{j}\right)^{2}-\left(\sigma_{j}\right)^{2}\right),$
where $D$ is the latent space dimension.
VQ-VAE assumes a uniform prior over the total $K$ codewords in the learnable codebook for the deterministic posterior, thus present a KL divergence as follows:
$\mathcal{L}_{\mathrm{kl}}\left(q_{\phi}(\mathbf{z})\|p(\mathbf{z})\right).$ (9)
[…]
$=-(-\log K)$
$=\log K$
A.1 Comparison to VQ
We list the advantages of SoftVQ in Tab. 5 with empirical support, and will add in the revision.
Table 5: Advantages of SoftVQ over VQ.
Advantages VQ SoftVQ Empirical Support
Grad. Broken yes no
No straight-through trick. SoftVQ-S-64 has a 1.41 rFID and 13.35 gFID.
Codebook Loss yes no
Not necessary. Adding codebook loss in SoftVQ-S-64
results in an 1.34 rFID of of but a worse gFID of 15.15.
Commit Loss yes no
Not necessary. Adding commit loss in SoftVQ-S-64
results in an rFID of 1.33 of but a worse gFID of 14.35.
Comp. Ratio low high
SoftVQ has much lower rFID$<$1.0 comapred to TiTok in Tab. 1.
Rep. Align. difficult easy
More discriminative features in main Fig. 4
While representation alignment can also applied to VQ, it fails to learn discriminative features with high compression ratio, mainly due to its broken gradient. Compared to SoftVQ, whose encoder and codebook’s parameters are directly learned by the alignment loss, the gradient of alignment loss on VQ will first be straight-through from decoder input to encoder output, and then the codebook is updated according to the codebook loss, resulting in indirect learning of discriminative features.
We provide an additional codewords visualization in Fig. 5, where SoftVQ learns code embeddings uniformly across the entire distribution.
Figure 5: Codewords visualization
Appendix B Additional Details of SoftVQ-VAE and its Variants
In this section, we present more details on the posterior of SoftVQ-VAE, its variant GMMVQ-VAE with the latent space as a GMM model, and the compatibility of SoftVQ-VAE with improvement techniques of VQVAE.
B.1 Posterior of SoftVQ-VAE
In SoftVQ, we similarly assume that the prior is a uniform distribution over the $K$ learnable codewords as in VQ, except for we have the posterior as the softmax probability:
$\mathcal{L}_{\mathrm{kl}}\left(q_{\phi}(\mathbf{z})\|p(\mathbf{z})\right).$ (10)
[…]
$=H(q_{\phi}(\mathbf{z}))-H(q_{\phi}(\mathbf{z}),p(\mathbf{z})),$
where $H(q_{\phi}(\mathbf{z})$ is the entropy for $H(q_{\phi}(\mathbf{z}),p(\mathbf{z}))$ is the cross-entropy between $q_{\phi}(\mathbf{z})$ and the uniform prior $p(\mathbf{z})\sim\mathcal{U}(0,K)$. In practice, $H(\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}[q_{\phi}(\mathbf{z})],p(\mathbf{z}))$ we compute the $\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}[q_{\phi}(\mathbf{z})]$ instead, where $\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}[q_{\phi}(\mathbf{z})]$ is computed on the averaged batch data during training.
B.2 GMMVQ-VAE
As discussed in the main paper, the latent space of SoftVQ can be interpreted as soft K-Means, and it can be further extended to a latent space of Gaussian Mixture Model. We present more details of this extension here.
In GMMVQ-VAE, the encoder predicts two things: the embedding $\hat{\mathbf{z}}$ and the weights of the Gaussian component $\omega(\hat{\mathbf{z}})$. We can then formulate the posterior as:
posterior: $\begin{split}&\textit{posterior: }q_{\phi}(\mathbf{z}|\mathbf{x})=\operatorname*{Softmax}\left(-\omega(\hat{\mathbf{z}})\|\hat{\mathbf{z}}-\mathcal{C}\|_{2}\right)\\ &\textit{latent: }\mathbf{z}=q_{\phi}(\mathbf{z}|\mathbf{x})\mathcal{C}\\ &\textit{kl: }\mathcal{L}_{\mathrm{kl}}=H(q_{\phi}(\mathbf{z}|\mathbf{x}))-H\left(\mathbb{E}_{\mathbf{x}\sim p(\mathbf{x})}q_{\phi}(\mathbf{z}|\mathbf{x})\right).\\ \end{split}$ (11)
The difference between SoftVQ and GMMVQ in our formulation is the way in computing the posterior, i.e., using fixed temperature parameter versus learning the data-dependent temperature parameters by the encoder. Note that we still maintain the codebook and its codewords entries directly for the decoder input. It is possible to make the latent space formally a GMM, by treating the codewords as Gaussian means, and formulating the decoder input with the re-parametrization trick. However, we find that this formulation with re-parametrization will hinder the learning of the latent space. Thus, we adopted the simpler design to use the codebook directly for reconstruction, assuming fixed variance in the Gaussian components of GMM.
B.3 Compatibility of SoftVQ-VAE
Figure 6: Illustration of product quantization (left) and residual quantization (right) with SoftVQ.
Since SoftVQ-VAE maintains the learnable codebook, previous techniques to improve VQ-VAE are also compatible with it. Here, we follow ImageFolder [62] to show the combination of SoftVQ with product quantization [45] and residual quantization [55], as illustrated in Fig. 6.
Product Quantization (PQ) [45] divides the latent codes into $G$ groups, with each group having its own codebook:
$\begin{split}\mathbf{z}&=[\mathbf{z}^{(1)},\mathbf{z}^{(2)},...,\mathbf{z}^{(G)}]\\ \quad q_{\phi}(\mathbf{z}^{(g)}|\mathbf{x})&=\operatorname*{Softmax}(-\|\hat{\mathbf{z}}^{(g)}-\mathcal{C}^{(g)}\|_{2}/\tau).\end{split}$ (12)
PQ can effectively increase the actual codebook size.
Residual Quantization (RQ) [55] applies multiple layers of quantization to the residual errors of the encoder output:
$\begin{split}\mathbf{z}_{l}&=\mathbf{z}_{l-1}+\operatorname{SoftVQ}(\mathbf{r}_{l-1})\\ \mathbf{r}_{l}&=\mathbf{r}_{l-1}-\mathbf{z}_{l}\\ \mathbf{z}_{0}&=0\\ \mathbf{r}_{0}&=\hat{\mathbf{z}}\end{split}$ (13)
where $\mathbf{r}$ is the residual and $l$ is the layer index. RQ captures more fine-grained features and thus better reconstruction.
Appendix C Additional Implementation Details
In this section, we provide more implementation details of the tokenizer training, the downstream generative models training, and the linear probing evaluation.
C.1 Implementation Details of SoftVQ-VAE
[…]
Similarly to Tian et al. [99] and Li et al. [62], we found that the discriminator is very important for training the tokenizer. Instead of using the common PatchGAN and StyleGAN architecture, we adopted the frozen DINO-S [9, 79] discriminator with a similar architecture to StyleGAN [48, 49], as in [99]. In addition, we use DiffAug [121], consistency regularization [119], and LeCAM regularization [101] for discriminator training as in [99]. The loss weight of consistency regularization is set to 4.0 and LeCAM regularization is set to 0.001. We did not use the adaptive discriminator loss weight since it is too tricky to tune.
C.2 Implementation Details of DiT, SiT, and MAR
[…]
We train all of our 256$\times$256 generative models on 8 MI250 GPUs and 512$\times$512 models on 8 MI300 GPUs.
C.3 Implementation Details of Linear Probing
We use the setup similar to that used in MAE [35], DAE [14], and REPA [3] to perform the linear probing evaluation in the latent tokenizer space and intermediate features of trained SiT models. Specifically, we train a parameter-free batch normalization layer and a linear layer for 90 epochs with a batch size of 16,384. We use the Lamb optimizer with cosine decay learning rate scheduler where the initial learning rate is set to 0.003.
Appendix D Additional Results
In this section, we present our ablation study on the SoftVQ-VAE tokenizer, more results of the trained generative models including results from longer training, and more visualization of the reconstruction results from the tokenizer and the generative models.
D.1 Ablation Study
Table 6: Ablation study of SoftVQ-VAE
Variants
Tokenizer SoftVQ-S + PQ (G=2) + PQ (G=4) + RQ + R-PQ (G=4) GMMVQ-S
rFID 1.33 1.19 1.03 1.55 1.17 1.12
Codebook Size
$K$ 512 1024 2048 4096 8192 16384
rFID 1.82 1.58 1.35 1.14 1.03 1.23
Latent Size
$L$/$D$ 64/16 64/32 64/64 32/32 32/64 32/128
rFID 2.31 1.03 0.63 1.61 1.24 0.79
Softmax Temp.
$\tau$ 0.001 0.01 0.07 0.1 1.0 learnable
rFID 2.03 1.37 1.03 1.19 1.52 1.10
The full ablation results are shown in Tab. 6.
SoftVQ variants. We compare our baseline SoftVQ-S with several variants, including different quantization approaches. Adding product quantization (PQ) with group size G=2 and G=4 both improve the performance, reducing rFID from 1.33 to 1.19 and 1.03 respectively. Although residual quantization (RQ) alone increases rFID to 1.55, combining it with product quantization (R-PQ) yields better results with rFID of 1.17. The GMMVQ-S variant also shows competitive performance with an rFID of 1.12.
Codebook size. We investigate the impact of codebook size K ranging from 512 to 16384. The results show that larger codebooks generally lead to better performance, with rFID decreasing from 1.82 (K=512) to 1.03 (K=8192). However, further increasing K to 16384 slightly degrades performance with an rFID of 1.23, suggesting that an overly large codebook may harm model stability.
Latent size. The latent dimensions L/D significantly affect the model performance. Starting from a small latent size of 64/16, increasing the dimensions initially helps, with 64/32 achieving the best rFID of 1.03. While further enlargement leads to better reconstruction performance, as seen in the 64/64 configuration with an rFID of 0.63, we find that the larger dimension of the latent space makes the generative models more challenging to learn.
Softmax temperature. The softmax temperature $\tau$ controls the sharpness of the posterior probability. We observe that very low temperatures $\tau=$0.001 result in poor performance, while moderate values around 0.07 achieve optimal results. Making $\tau$ learnable produces competitive performance, suggesting that adaptive temperature could be a practical choice.
D.2 Main Results
Table 7: System-level comparison on ImageNet 256$\times$256 conditional generation. We compare with both diffusion-based models and auto-regressive models with different types of tokenizers. Grey denotes SoftVQ. Our DiT/SiT results are reported with a total training of 3M iterations (compared to 4M of SiT/XL-2 + REPA and 7M for SiT/XL2 and DiT-XL/2), and MAR results are reported with a total training of 500 epochs with smaller learning rate (compared to 800 epochs of MAR-H). indicates results with DPM-Solver and 50 steps.
Gen. Model Tok. Model # Tokens $\downarrow$ Tok. rFID $\downarrow$ Gflops $\downarrow$ Throughput (imgs/sec) $\uparrow$ w/o CFG w/ CFG
gFID $\downarrow$ IS $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ gFID $\downarrow$ IS $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$
Auto-regressive Models
Taming-Trans. [23] VQ 256 7.94 - - 5.20 290.3 - - - - - -
RQ-Trans. [55] RQ 256 3.20 908.91 7.85 3.80 323.7 - - - - - -
MaskGIT [10] VQ 256 2.28 - - 6.18 182.1 0.80 0.51 - - - -
MAGE [59] VQ 256 - - - 6.93 195.8 - - - - - -
LlamaGen-3B [98] VQ 256 2.08 781.56 2.90 - - - - 3.06 279.7 0.80 0.58
TiTok-S-128 [115] VQ 128 1.61 33.03 6.50 - - - - 1.97 281.8 - -
MAR-H [60] KL 256 1.22 145.08 0.12 2.35 227.8 0.79 0.62 1.55 303.7 0.81 0.62
SoftVQ-L 32 0.61 67.93 2.19 3.83 211.2 0.77 0.61 2.54 273.6 0.78 0.61
MAR-H SoftVQ-BL 64 0.65 86.55 0.89 2.81 218.3 0.78 0.62 1.93 289.4 0.80 0.61
Diffusion-based Models
LDM-4 [88] KL 4096 0.27 157.92 0.37 10.56 103.5 0.71 0.62 3.60 247.7 0.87 0.48
U-ViT-H/2 [4] KL 1024 0.62 128.89 0.98 - - - - 2.29 263.9 0.82 0.57
MDTv2-XL/2 [29] 125.43 0.59 5.06 155.6 0.72 0.66 1.58 314.7 0.79 0.65
DiT-XL/2 [80] 80.73 0.51 9.62 121.5 0.67 0.67 2.27 278.2 0.83 0.53
SiT-XL/2 [72] 81.92 0.54 8.30 131.7 0.68 0.67 2.06 270.3 0.82 0.59
+ REPA [116] 5.90 157.8 0.70 0.69 1.42 305.7 0.80 0.65
SoftVQ-B 0.89 8.94 9.83 113.8 0.70 0.61 3.91 264.2 0.81 0.54
SoftVQ-BL 0.68 8.81 9.22 115.8 0.71 0.61 3.78 266.7 0.82 0.54
DiT-XL SoftVQ-L 32 0.74 14.52 8.74 9.07 117.2 0.71 0.61 3.69 270.4 0.83 0.53
SoftVQ-B 0.88 4.70 6.62 129.2 0.75 0.62 3.29 262.5 0.84 0.54
SoftVQ-BL 0.65 4.59 6.53 131.9 0.75 0.62 3.11 268.3 0.84 0.56
DiT-XL SoftVQ-L 64 0.61 28.81 4.51 5.83 141.3 0.75 0.62 2.93 268.5 0.84 0.55
SoftVQ-B 0.89 10.28 7.99 129.3 0.70 0.63 2.51 301.3 0.76 0.62
SoftVQ-BL 0.68 10.12 7.67 133.9 0.70 0.63 2.44 308.8 0.76 0.63
SiT-XL SoftVQ-L 32 0.74 14.73 10.11 7.59 137.4 0.71 0.63 2.44 310.6 0.77 0.63
SoftVQ-B 0.88 5.51 5.98 138.0 0.74 0.64 1.78 279.0 0.80 0.63
SoftVQ-BL 0.65 5.42 5.80 143.5 0.74 0.64 1.88 287.9 0.80 0.63
SiT-XL SoftVQ-L 64 0.61 29.23 5.39 5.35 151.2 0.74 0.64 1.86 293.6 0.81 0.63
Table 8: System-level comparison on ImageNet 512$\times$512 conditional generation. We compare with both diffusion-based models and auto-regressive models with different types of tokenizers. Grey denotes SoftVQ. indicates results with DPM-Solver and 50 steps.
Gen. Model Tok. Model # Tokens $\downarrow$ Tok. rFID $\downarrow$ Gflops $\downarrow$ Throughput (imgs/sec) $\uparrow$ w/o CFG w/ CFG
gFID $\downarrow$ IS $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$ gFID $\downarrow$ IS $\uparrow$ Prec. $\uparrow$ Recall $\uparrow$
Generative Adversarial Models
BigGAN [10] - - - - - - - - - 8.43 177.9 - -
StyleGAN-XL [48] - - - - - - - - - 2.41 267.7 - -
Auto-regressive Models
MaskGIT [10] VQ 1024 1.97 - - 7.32 156.0 - - - - -
TiTok-B [115] VQ 128 1.52 - - - - - - 2.13 261.2 - -
MAR-H SoftVQ-BL 64 0.71 86.55 1.50 8.21 152.9 0.69 0.59 3.42 261.8 0.77 0.61
Diffusion-based Models
ADM [16] - - - - - 23.24 58.06 - - 3.85 221.7 0.84 0.53
U-ViT-H/4 [4] KL 4096 0.62 128.92 0.58 - - - - 4.05 263.8 0.84 0.48
DiT-XL/2 [80] 373.34 0.10 9.62 121.5 - - 3.04 240.8 0.84 0.54
SiT-XL/2 [72] 373.32 0.10 - - - - 2.62 252.2 0.84 0.57
DiT-XL [11] AE 256 0.22 80.75 1.02 9.56 - - - 2.84 - - -
UViT-H [11] 128.90 6.14 9.83 - - - 2.53 - - -
UViT-H [11] AE 64 0.22 32.99 15.23 12.26 - - - 2.66 - - -
UViT-2B [11] 104.18 9.44 6.50 - - - 2.25 - - -
SoftVQ-L 32 0.64 14.73 10.12 10.17 119.2 0.65 0.59 4.23 218.0 0.83 0.52
SiT-XL SoftVQ-BL 64 0.71 29.23 5.38 7.96 133.9 0.73 0.63 2.21 290.5 0.85 0.59
We report the full results, including precision and recall, for the ImageNet 256$\times$256 benchmark in Tab. 7 and 512$\times$512 benchmark in Tab. 8, respectively. All main results reported are evaluated on SiT-XL and DiT-XL models trained for 3M iterations, and MAR models trained for 500 epochs. Noteworthy is that our models achieve performance comparable to state-of-the-art systems. More importantly, on 512$\times$512 generation, SoftVQ with 32 tokens provides a 100x speedup on the inference throughput.
An interesting finding is that, even though our models in general present the best conditional gFID without using CFG, their performance improvement with CFG becomes smaller, compared to SiT-XL/2 + REPA [116]. We leave the investigation of the reasons behind this observation for our future work.
We also provide results of longer training, i.e., 4M iterations, and results along the training in Tab. 9. One can observe that the performance of our models does not saturate at 3M iterations of training, and longer training may provide further performance improvement.
Table 9: Generation performance over training of DiT-XL and SiT-XL trained on SoftVQ-L with 32 and 64 tokens.
Model Training Iter. w/o CFG w/ CFG
FID IS Prec. Recall FID IS Prec. Recall
DiT/XL SoftVQ-L 32 3M 9.07 117.2 0.71 0.61 3.69 270.4 0.83 0.53
4M 8.54 124.0 0.72 0.62 3.58 281.9 0.82 0.53
DiT/XL SoftVQ-L 64 3M 5.83 141.3 0.75 0.62 2.93 268.5 0.84 0.55
4M 5.60 144.5 0.76 0.63 2.84 270.1 0.85 0.54
SiT/XL SoftVQ-L 32 400K 16.49 79.1 0.65 0.61 4.61 281.2 0.86 0.47
1M 10.30 109.0 0.69 0.63 2.85 284.0 0.76 0.61
2M 8.52 122.8 0.71 0.64 2.51 296.3 0.77 0.63
3M 7.59 137.4 0.71 0.63 2.44 310.6 0.77 0.63
4M 7.32 138.6 0.72 0.63 2.38 311.5 0.77 0.63
SiT/XL SoftVQ-L 64 400K 10.03 103.4 0.71 0.62 3.12 236.3 0.77 0.61
1M 7.14 125.4 0.72 0.64 2.09 254.9 0.79 0.62
2M 5.88 140.6 0.73 0.64 1.89 284.4 0.80 0.63
3M 5.35 151.2 0.74 0.64 1.86 293.6 0.81 0.63
4M 5.21 158.5 0.75 0.64 1.79 297.8 0.81 0.64
D.3 More Comparison to TiTok
We provide a comparison between TiTok and SoftVQ in Tab. 10, including token numbers, model parameters, rFID of IN-1K and MSCOCO, linear probing (LP) accuracy, gFID and IS of SiT-L. We show that SoftVQ significantly outperforms TiTok with VQ and KL. Also, we need to note that while the model architectures are similar, training is different. TiTok is trained using 2-stage decoders with a pre-trained tokenizer, whereas our method is trained end-to-end with a single decoder. Also note that, although soft assignment of codewords may not be a new idea, our work is indeed the first to explore it with continuous tokenizer that achieves high compression ratio.
Table 10: Comparison between TiTok and SoftVQ.
Tokenizer # Param. rFID (IN-1K) rFID (COCO) LP (IN-1K) gFID IS
TiTok-B-64 176 1.70 11.20 39.2 25.12 53.1
TiTok-BL-KL-64 389 1.25 8.94 11.4 23.35 52.7
TiTok-BL-VQ-64 390 2.60 11.24 9.3 19.23 61.8
SoftVQ-B-64 173 0.89 5.16 42.3 10.13 103.4
TiTok-L-32 614 2.21 14.41 48.9 26.84 49.5
TiTok-LL-KL-32 614 1.61 10.45 12.3 24.65 51.97
SoftVQ-L-32 608 0.61 5.51 59.4 9.47 107.2
D.4 Reconstruction Visualization
We present a visualization of the SoftVQ-L reconstruction results with 32 and 64 tokens in Fig. 7 and Fig. 8, respectively.
Figure 7: Reconstruction results of SoftVQ-L with 32 tokens.
Figure 8: Reconstruction results of SoftVQ-L with 64 tokens.
D.5 Generation Visualization
More visualizations of SiT-XL trained on SoftVQ-L with 32 and 64 tokens are shown here.
Figure 9: Uncurated 256$\times$256 generation results of DiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = ”balloon” (417).
Figure 10: Uncurated 256$\times$256 generation results of DiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “baseball” (429).
Figure 11: Uncurated 256$\times$256 generation results of DiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “dog sled” (537).
Figure 12: Uncurated 256$\times$256 generation results of DiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “space shuttle” (812).
Figure 13: Uncurated 256$\times$256 generation results of DiT-XL with SoftVQ-L 32 tokens. We use CFG with 4.0. Class label = “panda” (388).
Figure 14: Uncurated 256$\times$256 generation results of DiT-XL with SoftVQ-L 32 tokens. We use CFG with 4.0. Class label = “geyser” (974).
Figure 15: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “golden retriever” (207).
Figure 16: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “arctic wolf” (270).
Figure 17: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “lion” (291).
Figure 18: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “loggerhead sea turtle” (33).
Figure 19: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “fire truck” (555).
Figure 20: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “laptop” (620).
Figure 21: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “macaw” (88).
Figure 22: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “sulphur-crested cockatoo” (89).
Figure 23: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “ice cream” (928).
Figure 24: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 4.0. Class label = “cheeseburger” (933).
Figure 25: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 32 tokens. We use CFG with 4.0. Class label = “coral reef” (973).
Figure 26: Uncurated 256$\times$256 generation results of SiT-XL with SoftVQ-L 32 tokens. We use CFG with 4.0. Class label = “lake shore” (975).
Figure 27: Uncurated 512$\times$512 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 2.0. Class label = “cliff drop-off” (972).
Figure 28: Uncurated 512$\times$512 generation results of SiT-XL with SoftVQ-L 64 tokens. We use CFG with 2.0. Class label = “volcano” (980).